# Performance Improvement of Deorbitalized Exchange-Correlation Functionals

### v1

@@ II.3 Kinetic and exchange potentials
The smoothing problems that arise in the deorbitalizers are well illustrated by a plot of the kinetic potential $v_{\theta}$ (Eq. (19)) for the Hydrogen atom for various Laplacian-dependent KED functionals. See Fig. 2.
@@
Most of the approximate KED functionals deliver potentials that are reasonably close to the exact value (zero) except in the vicinity of 2.5 to 3.0 a.u. There violent fluctuations in the potential are evident. Those can be traced to a spurious diagnosis of density behavior. At high densities, the KED functionals diagnose hydrogen to be an iso-orbital system, which is correct. But at low density they treat it like a slowly varying electron gas. The fluctuations occur for the values of $r$ at which the transition between the two limits occurs. The problem is worst for PC which suffers oscillations larger than 100 hartrees. The problem is somewhat less severe for PCopt and RPP with oscillations in the tens of hartrees. Only the CR model Cancio and Redd (2017) and its extension to the RPP form, SRPP2, show more reasonable variations, on the order of a few hartrees. In fact, the size of those variations is correlated closely (with the exception of PCopt) with the abruptness of the function used to describe the transition (from iso-orbital to slowly varying electron gas) in the enhancement factor, as seen in Fig. 1. As a caution, we note that the plot shows potentials evaluated with the exact ground-state density for hydrogen. The unreasonable oscillatory behavior almost certainly would be worse for self-consistent densities from the associated potentials.
@@
Figures 3 and 4 show plots of the local part of the exchange potential (Eq. 18) for the parent meta-GGA functional (r2SCAN) and the full X potential for three deorbitalized versions of it (PCopt, RPP, and SRPP). The systems are the H atom (at two different length scales) and the Si atom. (Aside: r2SCAN deorbitalized with the RPP form has the formal functional name OFR2, but for simplicity in making comparisons with other deorbitalization options, we will continue to refer to it as RPP.) Note that the X potential in Fig. 3 is multiplied by the radial coordinate to bring out features nearer the valence edge where most variation in potentials occurs. For r2SCAN, the contribution from the functional derivative with respect to the kinetic energy density is omitted, as it is incorporated in the physics in the gKS scheme via Eq. 4. These potentials are computed using a modified version of the Python code densities Mejía-Rodriguez (2023) which employs analytic spherical Hartree-Fock densities. They were evaluated on a double-exponential grid with 300 to 10,000 radial points.
Though an unequivocal comparison of the gKS local part and ordinary KS potentials is not possible, it is clear that deorbitalization of r2SCAN introduces unphysical oscillations. In particular, we see large noise in the region $r\approx 3$ au for the H atom, where the Pauli potential of Fig. 2 is seen to fail. However, the SRPP deorbitalization of r2SCAN results in modestly less noise in the potential compared to the PCopt or RPP deorbitalizations. A similar trend occurs for Si, with SRPP slightly smoother than RPP and both smoother than PCopt. Interestingly, the RPP and SRPP produce potentials noticeably smoother overall than the local part of the parent metaGGA, despite use of the density Laplacian.
The noticeable bump in the r2SCAN H atom potential at large $r$ [Fig. 3(a)] may be a remnant of a major issue in the original SCAN. In a region of exponentially decaying density, $v_{x}$ for SCAN goes to a large positive value rather than zero Bartók and Yates (2019); Furness et al. (2020). That unphysical behavior forces the use of extremely fine grids and leads to difficulties in obtaining pseudopotentials.Holzwarth et al. (2022) (Note that the exact KS exchange potential for the H atom must cancel the self-Hartree potential, hence $rv_{x}$ must go asymptotically to 1-1, which no potential considered here does properly.)
The short-range behavior for any atom is exemplified by the H atom case. All potentials for it fail by diverging at extremely small distances, rather than trending to a small negative value. van Leeuwen and Baerends (1994). Overall $v_{x}$ from PCopt and SRPP are rather close. Unlike $v_{x}$ from RPP, they are roughly similar quantitatively to the parent $v_{x}$.
@@ II.4 Noise problem quantification
This technique has been applied previously to minimize the Laplacian term of the overall exchange-correlation potential. See Ref. Cancio et al., 2012 for details. Our choice to focus on the kinetic potential reflects the fact that the Laplacian term of the X potential actually generates two plausibly problematic contributions. One, $\nabla^{2}(\partial e_{\scriptscriptstyle\rm XC}/\partial\tau)$, arises from the functional dependence of $e_{\scriptscriptstyle\rm XC}$ on $\tau$. The other, $\nabla^{2}(\partial\tau/\partial\nabla^{2}n)$, arises from the dependence of the deorbitalization of $\tau$ on $\nabla^{2}n$. The effects of the form of $e_{\scriptscriptstyle\rm XC}(\tau)$ upon performance characteristics of both conventional and deorbitalized meta-GGAs has been the subject of recent attention. Bartók and Yates (2019); Furness et al. (2020); Kaplan and Perdew (2022) However, comparison of Fig. 2 to Fig. 3 shows that the location of major error in the hydrogen X potential (the bump in deorbitalized potentials around 2 to 3 au) coincides with that of dramatic noise in the Pauli potential. This implicates, strongly, the second term as a culprit in generating X potential noise, hence that term is a focus of concern. A second benefit of this choice is that it focuses on an aspect of the potential that is independent of the choice of XC functional being deorbitalized.
Minimization of oscillations in that term on a useful test density or densities would have the potential of minimizing its impact on numerical performance in general. Perhaps that should be be a design goal for a deorbitalized functional. We will explore that idea in a followup to the present study. For now, to get an efficient, useful estimate of the impact of deorbitalizer-generated noise, we simply evaluate Eq. (41) for the H atom as calculated from a given KED functional. The crucial point is that the exact Pauli potential (Eq. (19) for this case is zero.
The consequences of that evaluation are shown, along with the kinetic energy obtained by integrating $\tau_{\scriptscriptstyle\rm OF}$, in Table 1 with the deorbitalizers PC, PCopt, CR, RPP, SRPP, and SRPP2. There is a close correlation between the smoothness of transition between slowly-varying and iso-orbital limits exhibited by the enhancement factors seen in Fig. 1 and the quantified measure $I^{deorb}$. The switching behavior of PC clearly is the most abrupt. Thus it has an enormous $I^{deorb}$ value. PCopt  and RPP do much better, with a reduction in $I^{deorb}$ by a factor of five and an order of magnitude respectively. Clearly, the functionals with the CR switch (CR, SRPP, SRPP2) are the best. That pattern is manifest to a considerable extent in the corresponding kinetic potentials (Figs. 2) and to a lesser but still significant effect for X potentials, 3). The next issue, obviously, is to address to what extent (if any) this smoothing endeavor and its quantitative assessment has consequences in calculations on real systems.
@@ III Computational methods
For molecular system tests, calculations were done with the NWChem 7.0.2 code Aprà et al. (2020) using the def2-TZVPP basis set and xfine grid settings. Test comparisons are to heats of formation computed according to the now-standard procedure of Curtiss et al. Curtiss et al. (1997, 2001) for the 223 molecules of the G3X/99 test, to bond lengths for the T96-R test set Staroverov et al. (2003, 2004), and to harmonic vibrational frequencies using the T82-F test set Staroverov et al. (2003, 2004).
[…]
For timing calculations, the same methodology was used except that for the 55 solids Peng et al. (2016) test set, the calculations were redone at the experimental lattice constants. In that way only single-point energy evaluations were needed for timings. Calculations used a single node on the University of Florida HiperGator system (Gen. 3) with an AMD EPYC 75F3 32-core processor with 4GB of memory per core. They were executed in parallel across all 32 cores.
@@ IV Results
IV.1 Structural properties
Evidently the comparison of interest is primarily for r2SCAN versus its deorbitalized versions, r2SCAN-L, built with the PCopt deorbitalizer Mejía-Rodríguez and Trickey (2020), with RPP Kaplan and Perdew (2022), and with SRPP and SRPP2. An underlying issue, discussed in detail above, is that PCopt breaks some constraints that those other three preserve. This motivates further comparison, between PCopt and the three more constraint-compliant deorbitalizers.
Tables 2 and 3 compile mean errors (ME) and mean absolute deviations (MADs) for r2SCAN, and deorbitalized r2SCAN, for the four deorbitalizers, along with results from PBE Perdew et al. (1996) as a baseline. In addition to these standard measures of functional performance, we include the spread of errors as defined by the maximum (most positive) error minus the minimum (most negative.) This last measure is motivated by the wide variations in timing performance that we discuss below. That variability raises the question whether there might be similar variability in predictive performance.
@@ IV.1 Structural properties
For molecular systems, the results presented in Table 2 show that SRPP is a substantial improvement (40% reduction) over RPP for heat of formation MAD. SRPP2 also provides an improvement over RPP, albeit a significantly smaller one. On bond lengths and vibrational frequencies, the three (RPP, SRPP, SRPP2) perform about the same; none proves competitive with PCopt. All three RPP-derived deorbitalizations show a disappointing degradation of performance, especially for heats of formation and frequencies, compared to PCopt deorbitalization, and none can be considered a completely faithful deorbitalization of r2SCAN. Also note the near equality of the magnitude of MEs and MADs for SRPP2 and to a lesser extent, SRPP. This, and the fact that the MEs for both are negative, indicate that they tend to underbind consistently, leading to larger MADs.
@@
For solids, Table 3 shows quite different outcomes. SRPP, SRPP2, and RPP all have distinctively better MADs than PCopt for lattice constants, cohesive energies, and bulk moduli. SRPP and SRPP2 are a tiny bit better than RPP on lattice constant MAD, slightly worse on cohesive energy, and about 9% better on bulk modulus. The bulk modulus spread for SRPP and SRPP2 is better than for RPP as well and substantially better than that from PCopt .
@@
Overall, SRPP and SRPP2 preserve the formal properties of RPP and either improve on its performance or maintain it. RPP is already known to be better on metallic solids than PCopt, (and better in some measures than r2SCAN) but worse on molecules Kaplan and Perdew (2022). That behavior is confirmed by our results (again, see Tables 2 and 3). Among the three, SRPP and SRPP2 are slightly better than RPP for solids, while SRPP is the best of the three over all systems though not competitive with PCopt for molecules. These outcomes reinforce the finding by Kaplan and Perdew Kaplan and Perdew (2022) that re-introduction of compliance with constraints broken by PCopt actually can reduce the breadth of applicability of the deorbitalization.
@@
The error spread reduction provided by the RPP-derived models was unexpected. Those spreads are quite a bit better than the original r2SCAN for both cohesive energies and lattice constants and not much worse for bulk moduli. In that regard, PCopt shows some degradation. Similar results are found if we measure standard deviations. That is to say, deorbitalization can, in some cases, lead to a significant reduction in outliers compared to the gKS calculations with the parent metaGGA XC functional, an unexpected bonus. The inverse correlation of spread to our measure of potential smoothness is suggestive. If some of that spread is due to numerical instability rather than functional accuracy, these results would be a trade-off between removing numerical problems with gKS and introducing problems with the Laplacian. The relative absence of outliers leads us to speculate that the noise reduction in smoothed deorbitalized potentials may reduce the risk that geometry optimization procedures will discover spurious local minima.
IV.2 Timing results

### v2

@@ II.3 Kinetic and exchange potentials
The smoothing problems that arise in the deorbitalizers are well-illustrated by a plot of the kinetic potential $v_{\theta}=\delta T_{\theta}/\delta n$, Eq. (20), for various relevant Laplacian-dependent KED functionals for the Hydrogen atom, evaluated with the exact H atom density. See Fig. 2.
@@
Most of the approximate KED functionals deliver potentials that are reasonably close to the exact value (zero) except in the vicinity of 2.5 to 3.0 a.u. There violent fluctuations in the potential are evident. Those can be traced to a spurious diagnosis of density behavior. At high densities, the KED functionals diagnose hydrogen to be an iso-orbital system, which is correct. But at low density they treat it like a slowly varying electron gas. The fluctuations occur for the values of $r$ where the transition between the two behaviors occurs. The problem is worst for PC which suffers oscillations larger than 100 hartrees. The problematic behavior is somewhat less severe for PCopt and RPP with oscillations in the tens of hartrees. Only the CR model [25] and its extension to the RPP form, SRPP2, show more reasonable variations, on the order of a few hartrees. In fact, the size of those variations is correlated closely (with the exception of PCopt) with the abruptness of the function used to describe the transition (from iso-orbital to slowly varying electron gas) in the enhancement factor, as seen in Fig. 1. As a caution, we note that the plot shows potentials evaluated with the exact ground-state density for hydrogen. The unreasonable oscillatory behavior almost certainly would be worse for self-consistent densities from the associated potentials.
@@
Figures 3 and 4 show plots of the local part of the exchange potential (Eq. 18) for the parent meta-GGA functional (r2SCAN) and the full X potential for three deorbitalized versions of it (PCopt, RPP, and SRPP). The systems are the H atom (at two different length scales) and the Si atom. (Aside: r2SCAN deorbitalized with the RPP form has the formal functional name OFR2. As a generalizable name for any deorbitalized meta-GGA that is reasonably compatible with previous usage, e.g. M-RT, from here on we have adopted the portmanteau of meta-GGA, “-L" (for Laplacian dependence), and deorbitalizer name in parentheses. Our new deorbitalizations of r2SCAN thus are r2SCAN-L(SRPP) and (SRPP2). For simplicity in making comparisons between deorbitalization options, where context allows, frequently we will refer to the deorbitalizer name, not the full functional name.) Note that the X potential in Fig. 3 is multiplied by the radial coordinate to bring out features near the valence edge where most variation in potentials occurs. For r2SCAN, the contribution from the functional derivative with respect to the kinetic energy density is omitted, as it is incorporated in the physics of the gKS scheme via Eq. (4). These potentials are computed using a modified version of the Python code densities [38] which employs fixed analytic spherical densities obtained from self-consistent Hartree-Fock calculations. [39] They were evaluated on a double-exponential grid with 1000 radial points, with substantially more points for r22SCAN.
@@
First, we note a comparison with a prior calculation of the r2SCAN-L(RPP) (OFR2) X potential as a function of radial distance for the Hydrogen atom, Fig. 7 of Ref. 27. Even allowing for how the information is displayed, $rV_{\scriptscriptstyle\rm X}(r)$ on the $y$ axis here versus $V_{\scriptscriptstyle\rm X}(r)$ in the other paper and $r$ on a log scale here versus uniform in the other)], it is clear that the two noise profiles are profoundly different. Ours shows considerably less noise, assuming any sharp feature in the potential outside the cusp region to be noise induced. As shown in the Supplemental Information [40], we have analyzed the dependence of the numerically calculated X potential upon grid density. We find oscillations of magnitude similar to those shown in Ref. 27 when our grid has 300 points. Those vanish with finer grids and converge to the value shown here for a grid size greater than 700 points. We also verify that the two potentials, when plotted in the same fashion, are essentially identical up to noise. We do not know if the prior work included a noise sensitivity analysis, but doing so seems to be an important consideration when working with the density Laplacian.
Though an unequivocal comparison of the gKS local part and ordinary KS potentials is not possible, it is clear that deorbitalization of r2SCAN introduces unphysical oscillations. In particular, we see large noise in the region $r\approx 3$ au for the H atom, where the Pauli potential, Fig. 2, is seen to fail. However, the SRPP deorbitalization of r2SCAN [r2SCAN-L(SRPP)] results in modestly less noise in the potential compared to the PCopt or RPP deorbitalizations. A similar trend occurs for Si, with SRPP slightly smoother than RPP and both smoother than PCopt. Interestingly, the RPP and SRPP produce potentials noticeably smoother overall than the local part of the parent meta-GGA, despite use of the density Laplacian. (To be sure that this indeed is true, we needed to increase the resolution for the r2SCAN results to 10,000 points for H and 40,000 for Si.)
The noticeable bump in the r2SCAN H atom potential at large $r$ [Fig. 3(a)] may be a remnant of a major issue in the original SCAN. In a region of exponentially decaying density, $v_{x}$ for SCAN goes to a large positive value rather than zero [6, 7]. That unphysical behavior forces the use of extremely fine grids and leads to difficulties in obtaining pseudopotentials.[14] (Note that the exact KS exchange potential for the H atom must cancel the self-Hartree potential, hence $rv_{x}$ must go asymptotically to 1-1, which no potential considered here does properly.)
The short-range behavior for any atom is exemplified by the H atom case. All potentials for it fail by diverging at extremely small distances, rather than trending to a small negative value [41]. Overall $v_{x}$ from PCopt and SRPP are rather close. Unlike $v_{x}$ from RPP, they are roughly similar quantitatively to the parent $v_{x}$.
@@ II.4 Noise problem quantification
This technique has been applied previously to minimize the Laplacian term of the overall exchange-correlation potential – see Ref. 42 for details. Our choice to concentrate on the kinetic potential reflects the fact that the Laplacian term of the X potential actually generates two plausibly problematic contributions. One, $\nabla^{2}(\partial e_{\scriptscriptstyle\rm XC}/\partial\tau)$, arises from the functional dependence of $e_{\scriptscriptstyle\rm XC}$ on $\tau$. The other, $\nabla^{2}(\partial\tau/\partial\nabla^{2}n)$, arises from the dependence of the deorbitalization of $\tau$ on $\nabla^{2}n$. The effects of the form of $e_{\scriptscriptstyle\rm XC}(\tau)$ upon performance characteristics of both conventional and deorbitalized meta-GGAs has been the subject of recent attention. [6, 7, 27] However, comparison of Fig. 2 to Fig. 3 shows that the location of major error in the hydrogen X potential (the bump in deorbitalized potentials around 2 to 3 au) coincides with that of dramatic noise in the Pauli potential. This implicates, strongly, the second term as a culprit in generating X potential noise, hence that term is a focus of concern. A second benefit of this choice is to focus on an aspect of the potential that is independent of the choice of XC functional being deorbitalized.
Minimization of oscillations in that term on a useful test density or densities would have the potential of minimizing its impact on numerical performance in general. Perhaps that should be a design goal for a deorbitalized functional. We will explore that idea in a follow-up to the present study. For now, to get an efficient, useful estimate of the impact of deorbitalizer-generated noise, we simply evaluate Eq. (41) for the H atom as calculated from a given KED functional. The crucial point is that the exact Pauli potential, Eq. (20), for this case is zero.
The consequences of that evaluation are shown, along with the kinetic energy obtained by integrating $\tau_{\scriptscriptstyle\rm OF}$, in Table 1 with the deorbitalizers PC, PCopt, CR, RPP, SRPP, and SRPP2. There is a close correlation between the smoothness of transition between slowly-varying and iso-orbital limits exhibited by the enhancement factors seen in Fig. 1 and the quantified measure $I^{deorb}$. The switching behavior of PC clearly is the most abrupt. Thus it has an enormous $I^{deorb}$ value. PCopt  and RPP do much better, with a reduction in $I^{deorb}$ by a factor of five and an order of magnitude respectively. Clearly, the functionals with the CR switch (CR, SRPP, SRPP2) are the best. That pattern is manifest to a considerable extent in the corresponding kinetic potentials (Fig. 2) and to a lesser but still significant extent for X potentials, Fig. 3.
The next issue, obviously, is to address to what extent (if any) this smoothing endeavor and its quantitative assessment has consequences for calculations on real systems.
@@ III Computational methods
For molecular system tests, calculations were done with the NWChem 7.0.2 code [43] using the def2-TZVPP basis set and xfine grid settings. Test comparisons are to heats of formation computed according to the now-standard procedure of Curtiss et al. [44, 45] for the 223 molecules of the G3X/99 test, to bond lengths for the T96-R test set [46, 47], and to harmonic vibrational frequencies using the T82-F test set [46, 47].
Solid structural properties were calculated with the VASP-5.4.4 code [48], following the same methodology and protocol as in Ref. 49.
In particular, we used projector augmented wave (PAW) methodology with PBE-based datasets, consistent with previous M-RT studies [18]. (Since PAWs for meta-GGAs are unavailable in VASP, use of PBE PAW datasets for meta-GGAS in VASP calculations is accepted practice.) For hexagonal close-packed structures the ideal $c/a$ ratio was used. For cohesive energies, isolated atom energies were calculated using a $14\times 15\times 16\AA$ unit cell and $\Gamma$-point sampling. Static-crystal lattice constants and cohesive energies are compared to published results for 55 solids [50] and, correspondingly, for bulk moduli of 44 solids [51]. In addition, band gaps of 21 insulators and semiconductors [52] were computed. Equilibrium lattice constants $a_{0}$ and bulk moduli $B_{0}$ at $T=0^{\circ}K$ were determined by calculating the total energy per unit cell at twelve points in the volume range $V_{0}\pm 10\%$, where $V_{0}$ is the calculated equilibrium unit cell volume, followed by fitting to the stabilized jellium equation of state (SJEOS) [53]. All the error values are from comparison with tabulated experimental values.
For timing calculations, the same methodology was used except that for the 55 solids [50] test set, the calculations were redone at the experimental lattice constants. In that way only single-point energy evaluations were needed for timings. Calculations used a single node on the University of Florida HiperGator system (Gen. 3) with an AMD EPYC 75F3 32-core processor with 4GB of memory per core. They were executed in parallel across all 32 cores.
@@ IV Results
IV.1 Atoms
An issue raised by moving from the RPP to SRPP is norm compliance. The free coefficients $c_{1}$, $c_{2}$, and $c_{3}$, are determined in the RPP by fitting to 1) the XC energy of noble gas atoms Ne, Ar, Kr, Xe; 2) “cluster" energies of jellium clusters; and 3) surface energies of jellium clusters. In developing the SRPP with the goal of smoothing the potential relative to RPP, we have not modified any aspect of the RPP except for the transition between the slowly varying and iso-orbital limits. The $c$ coefficients and other parameters in the slowly-varying form of RPP are left unchanged. Thus, it would be unsurprising if the norms fit by them were to be less well met in SRPP than in RPP.
We can test the effect that replacing the piecewise polynomial transition function [Eqs. (26) and (37)] in RPP with the smoother SRPP function [Eq. (30)] has on satisfaction of the “appropriate norm" of atomic exchange-correlation energies. We calculated self-consistent RPP and SRPP exchange-correlation energies for atoms up through $Z=18$ using NWChem, with the same basis sets as for molecules. As benchmark energies, we take highly accurate correlation energies from Ref. 54 and exact exchange calculation values using the optimized effective potential code OPMKS [55].
Overall, we find a mean error of about 2.1% for the RPP and 2.8% for SRPP, both systematically negative. This error decreases gradually in magnitude with system size. That decrease is expected since the XC energy of large-$Z$ atoms is increasingly well-characterized by the slowly-varying electron gas as $Z\to\infty$ [56], and each functional matches that limit as a norm. The discrepancy between RPP and SRPP for atoms between $Z=10$ and $18$ is about 0.5%, decreasing rapidly with $Z$. This rapid decrease is consistent with the decrease in importance of the cusp region and the exponential decay of the density far from the nucleus, the two regions in which the transition from gradient expansion to von Weizsäcker behavior is most important. See the Supplemental Information for more details.
We lack the capacity to test compliance with the jellium norms, but expect similar results.
IV.2 Structural properties
Evidently the comparison of primary interest is for r2SCAN versus its deorbitalized versions, r2SCAN-L(PCopt) , with the PCopt deorbitalizer [19], as well as with r2SCAN-L(RPP) [27] and with r2SCAN-L(SRPP) and (SRPP2). An underlying issue, discussed in detail above, is that PCopt breaks some constraints that the other three preserve. That distinction motivates further comparison, namely between PCopt and the three more constraint-compliant deorbitalizers.
Tables 2 and 3 compile mean errors (ME) and mean absolute deviations (MADs) for r2SCAN, and deorbitalized r2SCAN, for the four deorbitalizers, along with results from PBE [57] as a baseline. In addition to those standard measures of functional performance, we include the spread of errors as defined by the maximum (most positive) error minus the minimum (most negative.) This last measure is motivated by the wide variations in timing performance that we discuss below. That variability raises the question whether there might be similar variability in predictive performance.
@@ IV.2 Structural properties
For molecular systems, the results presented in Table 2 show that SRPP is a substantial improvement (40% reduction) over RPP for heat of formation MAD. SRPP2 also provides an improvement over RPP, albeit a significantly smaller one. On bond lengths and vibrational frequencies, the three (RPP, SRPP, SRPP2) perform about the same. None proves competitive with PCopt. All three RPP-derived deorbitalizations show a disappointing degradation of performance, especially for heats of formation and frequencies, compared to PCopt deorbitalization, and none can be considered a completely faithful deorbitalization of r2SCAN. Also note the near equality of the magnitude of MEs and MADs for SRPP2 and to a lesser extent, SRPP. This, and the fact that the MEs for both are negative, indicate that they tend to underbind consistently, leading to larger MADs.
@@
For solids, Table 3 shows quite different outcomes. SRPP, SRPP2, and RPP all have distinctly better MADs than PCopt for lattice constants, cohesive energies, and bulk moduli. SRPP and SRPP2 are a tiny bit better than RPP on lattice constant MAD, slightly worse (3%) on cohesive energy, and about 9% better on bulk modulus. The bulk modulus spread for SRPP and SRPP2 is better than for RPP as well and substantially better than that from PCopt .
@@
Overall, SRPP and SRPP2 preserve the formal properties of RPP and either improve on its performance or maintain it. RPP is already known to be better on metallic solids than PCopt, (and better in some measures than r2SCAN) but worse on molecules [27]. That behavior is confirmed by our results (again, see Tables 2 and 3). Among the three, SRPP and SRPP2 are slightly better than RPP for solids, while SRPP is the best of the three over all systems though not competitive with PCopt for molecules. These outcomes reinforce the finding by Kaplan and Perdew [27] that re-introduction of compliance with constraints broken by PCopt actually can reduce the breadth of applicability of the deorbitalization.
@@
The error spread reduction provided by the RPP-derived models was unexpected. Those spreads are quite a bit better than the original r2SCAN for both cohesive energies and lattice constants and not much worse for bulk moduli. In that regard, PCopt shows some degradation. Similar results are found if we measure standard deviations. That is to say, deorbitalization can, in some cases, lead to a significant reduction in outliers compared to the gKS calculations with the parent meta-GGA XC functional, an unexpected bonus. The inverse correlation of spread to our measure of potential smoothness is suggestive. If some of that spread is due to numerical instability rather than functional accuracy, these results would be a trade-off between removing numerical problems with gKS and introducing problems with the Laplacian. The relative absence of outliers leads us to speculate that the noise reduction in smoothed deorbitalized potentials may reduce the risk that geometry optimization procedures will discover spurious local minima.
IV.3 Timing results